OpenStudy (anonymous):

Dusty planted a rectangular garden that had a perimeter of 36 yards. The length of the garden is 6 feet longer than twice the width. What is the width of the garden in feet

OpenStudy (anonymous):

2(W+6) +2W=36 2W + 12 +2W = 36 4W = 24 W = 6 L = W+6 = 6+6 = 12

OpenStudy (mertsj):